(a) A piston at 5 atm contains a gas that occupies a volume of 3.5 L. What pressure would have to be placed on the piston to force the volume to adjust to 1.05 L?
_________atm

(b) A piston at 30.0°C contains a gas that occupies a volume of 2.5 L. To what temperature would the gas have to be heated to increase the volume to 6.2 L at constant pressure?
________°C

(c) A piston at 2145 torr contains a gas that occupies a volume of 4.7 L. What pressure would have to be placed on the piston to force the volume to adjust to 1.54 L?
_________ atm
Answer #1

(a) Formula: P1V1 = P2V2 (According to Boyle's law)

Where P1 and P2 are the initial and final pressures, respectively; V1 and V2 are the initial and final volumes, respectively.

i.e. 5 atm * 3.5 L = P2 * 1.05 L

i.e. P2 = 16.7 atm

(b) Formula: V1/T1 = V2/T2 (According to Charle's law)

Where V1 and V2 are the initial and final volumes, respectively; T1 and T2 are the initial and final temperatures, respectively.

i.e. 2.5 L/(30+273.15) K = 6.2 L/T2

i.e. T2 = 751.812 K = 751.812 - 273.15 = 478.662 oC

(c) Formula: P1V1 = P2V2 (According to Boyle's law)

i.e. 2145 torr * 4.7 L = P2 * 1.54 L

i.e. P2 = 6546 torr

1 atm = 760 torr

Thereforer, P2 = 6546/760 = 8.6 atm